# Joint Discrete Precoding and RIS Optimization for RIS-Assisted MU-MIMO Communication Systems

### v1

@@ IV RIS with Finite Phase-Shift Resolution
where $\mathbf{A}$ and $\mathbf{t}$ are defined in (31) and (32), respectively. One way to solve problem (38), which is prevalent in the prior literature, is to first solve the problem for the case of a continuous RIS configuration and then map each obtained reflection coefficient to the nearest feasible point in the set $\mathcal{D}$ [21, 24, 25]. With this approach, suboptimal discrete RIS reflection coefficients are attained as
@@
where $\overline{\boldsymbol{\theta}}$ is optimized using the method described in Section III-B. This is a low-complexity approach for finding the discrete RIS configuration; however, it is far from optimal because the reflection coefficients are selected independently and the quantization errors pile up, which is particularly troublesome in multi-user setups where the errors lead to inter-user interference.
@@
The triangular structure of $\mathbf{R}_{\theta}$ allows us to leverage the SESD algorithm for solving problem (41). Note that here, we cannot reformulate the SESD problem as an equivalent real-valued form similar to the one in (22) by separating the real and imaginary parts of $\boldsymbol{\theta}$ because the real and imaginary parts of $\boldsymbol{\theta}$ are intertwined and cannot be separately optimized. Instead of searching over a line of equally-spaced integers in the precoding optimization problem, we have to search over the equally-spaced angles on the unit circle. Therefore, the inputs of Algorithm 1 for optimizing the RIS configuration will be $\mathbf{R}_{\theta}$, $\mathbf{d}_{\theta}$, $\mathcal{D}$, and $N$.
@@ IV-A Rank-Deficiency Problem in RIS Configuration Optimization
${\underset{{\boldsymbol{\theta}\in\mathcal{D}^{N}}}{\mathrm{minimize}}}\,\,\boldsymbol{\theta}^{\mbox{\tiny$\mathrm{H}$}}(\mathbf{A}+\mathbf{I}_{N})\boldsymbol{\theta}-\mathbf{t}^{\mbox{\tiny$\mathrm{H}$}}\boldsymbol{\theta}-\boldsymbol{\theta}^{\mbox{\tiny$\mathrm{H}$}}\mathbf{t}.$ (42)
@@
where $\mathbf{A}+\mathbf{I}_{M}=\tilde{\mathbf{R}}_{\theta}^{\mbox{\tiny$\mathrm{H}$}}\tilde{\mathbf{R}}_{\theta}$ and $\tilde{\mathbf{d}}_{\theta}=(\mathbf{t}^{\mbox{\tiny$\mathrm{H}$}}\tilde{\mathbf{R}}_{\theta}^{-1})^{\mbox{\tiny$\mathrm{H}$}}$. Since all the diagonal elements of $\tilde{\mathbf{R}}_{\theta}$ are non-zero, the SESD technique can be applied to solve problem (43).
@@ VI Numerical Results and Discussion
    Coordinate descent: In each iteration of the WMMSE algorithm, a coordinate descent approach is utilized for optimizing the discrete vector. Specifically, in this approach, entries of the vector of interest are iteratively optimized such that in each iteration, one of the entries is optimized using an exhaustive search over the discrete set while other entries are kept fixed.
@@
    Nearest point: The optimal continuous solution is first found and then each entry of the optimized continuous vector is mapped to its nearest value in the quantization set (i.e., set of quantization alphabets or set of discrete reflection coefficients). In particular, in nearest point precoding, we use the optimal infinite-resolution precoding vectors as given in (18) in each iteration of the WMMSE algorithm and quantize the entries of the precoding vectors obtained after convergence. For nearest point RIS configuration, the alternating optimization approach presented in Section III-B is employed in each iteration of the WMMSE algorithm and the obtained configuration after convergence is quantized.
@@ VI-A Simulation Setup
$\mathbf{a}_{\mathrm{RIS}}(\varphi_{\mathrm{AoA}},\phi_{\mathrm{AoA}})=\Big{[}1,e^{j\psi_{\mathrm{H}}},\ldots,e^{j(N_{\mathrm{H}}\psi_{\mathrm{H}}+N_{\mathrm{V}}\psi_{\mathrm{V}})}\Big{]}^{\mbox{\tiny$\mathrm{T}$}},$ (54)
@@ VI-B Simulation Results
Fig. 5 investigates the benefits of using the RIS for assisting the communication between the BS and the UEs. The blue curves indicate the scenarios with optimized RIS phase shifts where the optimization is performed as discussed in Section III-B, while the orange curves correspond to cases with random phase shifts which can represent scattering from a rough surface. In the latter case, the RIS phase shifts are once randomized, and the same values are then used throughout the simulations. Therefore, the problem was reduced to optimizing the precoding vectors only. Two quantization levels for precoding is considered: $L=4$ and $L=8$. It is observed that using an RIS with optimized phase shifts can improve the system performance which highlights the gain that can be brought by RIS into the system.
@@
In Fig. 6, we evaluate the performance of the proposed SESD-based precoding by comparing it with nearest point precoding and coordinate descent precoding. In this simulation, continuous RIS configuration is considered and two quantization levels at the fronthaul are invesigated: $L=4$ and $L=8$. The ideal scenario with infinite-resolution precoding is also plotted which outperforms all the schemes with discrete precoding levels due to its higher interference suppression capability. We can see in Fig. 6 that the sum rate of the SESD-based precoding is much superior to those of nearest point precoding and coordinate descent precoding, and the gap between the our SESD-based design and the benchmarks increases with increasing the transmit power. That is because the benchmark schemes fail to efficiently suppress the interference among UEs since they only obtain suboptimal precoding vectors. On the contrary, the SESD-based precoding which finds the optimal discrete precoding vectors enjoys performance improvement when a larger transmit power is used at the BS. The sum rate will saturate at high SNRs also with the proposed method, but at a much higher level that is not visible in this figure.
@@
An important observation can be made by comparing Figs. 6 and 7. The gap between the SESD-based approach and the benchmarks is greater in Fig. 6 than in Fig. 7. Although an RIS with continuous phase shifts is considered in Fig. 6, the performance of the benchmark schemes that employ sub-optimal precoding vectors is very poor. When an optimal precoding design is considered in Fig. 7, the benchmarks demonstrate a better performance, although the RIS phase shifts are still sub-optimally found in these schemes. This suggests that the optimal precoder design is more important for improving performance as compared to the optimal RIS configuration design. The reason is that with the sub-optimal RIS configuration designs, only the phase shifts are affected as the reflection coefficients of the RIS are assumed to have unit modulus. On the other hand, a sub-optimal precoding design also affects the amplitude of the precoding entries, which has a large impact on the achievable sum rate.
@@ VI-C Discussion
VII Concluding Remarks
@@
We studied an RIS-assisted MU-MIMO downlink communication scenario in this paper. Since the precoding is computed in the cloud in future networks, the precoding vectors must be selected from a discrete set. Furthermore, a practical RIS typically has a low phase-shift resolution per element. We developed a novel framework for sum-rate maximization considering precoding vectors and RIS configurations adhering to the above-mentioned limitations. We devised efficient solutions based on the SESD algorithm for both precoding and RIS configuration design and compared them against the well-known coordinate descent and nearest point designs. Numerical results demonstrated the effectiveness of the proposed designs and revealed the noticeable performance gains of the presented schemes over the commonly used benchmarks. The SESD complexity grows exponentially with respect to the problem size, making it unsuitable for practically large RIS. Hence, we developed a novel low-complexity SESD-based algorithm that divides the discrete RIS configuration optimization problem into a number of sub-problems, each solved via SESD. The effectiveness of the proposed low-complexity algorithm has also been corroborated numerically. This low-complexity approach can be utilized to solve any mixed-integer least-squares problem where the large problem size inhibits the use of conventional SD-based algorithms.
The proposed SESD algorithms solve least-squares problems with discrete variables. Hence, there are good prospects that the proposed framework can also be applied to other RIS-related problems in future works.

### v2

@@ IV RIS with Finite Phase Shift Resolution
where $\mathbf{A}$ and $\mathbf{t}$ are defined in (31) and (32), respectively. One way to solve problem (38), which is prevalent in the prior literature, is to first solve the problem for the case of a continuous RIS configuration and then map each obtained reflection coefficient to the nearest feasible point in the set $\mathcal{D}$. With this approach, sub-optimal discrete RIS reflection coefficients are attained as
@@
where $\overline{\boldsymbol{\theta}}$ is obtained using the method described in Section III-B. This is a low-complexity approach for finding the discrete RIS configuration; however, it is far from optimal because the reflection coefficients are selected independently and the quantization errors pile up, which is particularly troublesome in multi-user setups where the errors lead to inter-user interference.
@@
The triangular structure of $\mathbf{R}_{\theta}$ allows us to leverage the SESD algorithm for solving problem (41). Note that here, we cannot reformulate the SESD problem as an equivalent real-valued form similar to the one in (22) by separating the real and imaginary parts of $\boldsymbol{\theta}$ because the real and imaginary parts of $\boldsymbol{\theta}$ are intertwined and cannot be separately optimized. Instead of searching over a line as in the precoding optimization problem, we have to search over the unit circle. Therefore, the inputs of Algorithm 1 for optimizing the RIS configuration will be $\mathbf{R}_{\theta}$, $\mathbf{d}_{\theta}$, $\mathcal{D}$, and $N$.
@@ IV-A Rank-Deficiency Problem in RIS Configuration Optimization
${\underset{{\boldsymbol{\theta}\in\mathcal{D}^{N}}}{\mathrm{minimize}}}\,\,\boldsymbol{\theta}^{\mbox{\tiny$\mathrm{H}$}}(\mathbf{A}+\alpha\mathbf{I}_{N})\boldsymbol{\theta}-\mathbf{t}^{\mbox{\tiny$\mathrm{H}$}}\boldsymbol{\theta}-\boldsymbol{\theta}^{\mbox{\tiny$\mathrm{H}$}}\mathbf{t}.$ (42)
@@
where $\mathbf{A}+\alpha\mathbf{I}_{N}=\tilde{\mathbf{R}}_{\theta}^{\mbox{\tiny$\mathrm{H}$}}\tilde{\mathbf{R}}_{\theta}$ and $\tilde{\mathbf{d}}_{\theta}=(\mathbf{t}^{\mbox{\tiny$\mathrm{H}$}}\tilde{\mathbf{R}}_{\theta}^{-1})^{\mbox{\tiny$\mathrm{H}$}}$. Since all the diagonal elements of $\tilde{\mathbf{R}}_{\theta}$ are non-zero, the SESD technique can be applied to solve problem (43).
@@ VI Numerical Results and Discussion
    Coordinate descent: In each iteration of the WMMSE algorithm, a coordinate descent approach is utilized for optimizing the discrete vector. Specifically, in this approach, entries of the vector of interest are iteratively optimized such that in each iteration, one of the entries is optimized using an exhaustive search over the discrete set while other entries are kept fixed. This approach has been used in [22, 23] for optimizing the discrete RIS configuration.
@@
    Nearest point: The optimal continuous solution is first found and then each entry of the optimized continuous vector is mapped to its nearest value in the discrete set (i.e., set of quantization alphabets or set of discrete reflection coefficients). In particular, in nearest point precoding, we use the optimal infinite-resolution precoding vectors as given in (18) in each iteration of the WMMSE algorithm and quantize the entries of the precoding vectors obtained after convergence. For nearest point RIS configuration, the alternating optimization approach presented in Section III-B is employed in each iteration of the WMMSE algorithm and the obtained configuration after convergence is quantized. References [21, 24, 25] have utilized this method for finding the discrete RIS configuration.
@@ VI-A Simulation Setup
$\mathbf{a}_{\mathrm{RIS}}(\varphi_{\mathrm{AoA}},\phi_{\mathrm{AoA}})=\big{[}1,e^{j\psi_{\mathrm{H}}},\ldots,e^{j\left((N_{\mathrm{H}}-1)\psi_{\mathrm{H}}+(N_{\mathrm{V}}-1)\psi_{\mathrm{V}}\right)}\big{]}^{\mbox{\tiny$\mathrm{T}$}},$ (54)
@@ VI-B Simulation Results
Fig. 5 investigates the benefits of using the RIS for assisting the communication between the BS and the UEs. The blue curves indicate the scenarios with optimized RIS phase shifts where the optimization is performed as discussed in Section III-B, while the orange curves correspond to cases with random phase shifts which can represent scattering from a rough surface. In the latter case, the RIS phase shifts are once randomized, and the same values are then used throughout the simulations. Therefore, the problem reduces to optimizing the precoding vectors only. Two quantization levels for precoding is considered: $L=4$ and $L=8$. It is observed that using an RIS with optimized phase shifts can improve the system performance which highlights the gain that can be brought by RIS into the system.
@@
In Fig. 6, we evaluate the performance of the proposed SESD-based precoding by comparing it with nearest point precoding and coordinate descent precoding. In this simulation, continuous RIS configuration is considered and two quantization levels at the fronthaul are invesigated: $L=4$ and $L=8$. The ideal scenario with infinite-resolution precoding is also plotted which outperforms all the schemes with discrete precoding levels due to its higher interference suppression capability. We can see in Fig. 6 that the sum rate of the SESD-based precoding is much superior to those of nearest point precoding and coordinate descent precoding, and the gap between the our SESD-based design and the benchmarks increases with increasing the transmit power. That is because the benchmark schemes fail to efficiently suppress the interference among UEs since they only obtain sub-optimal precoding vectors. On the contrary, the SESD-based precoding which finds the optimal discrete precoding vectors enjoys performance improvement when a larger transmit power is used at the BS. The sum rate will saturate at high SNRs also with the proposed method, but at a much higher level that is not visible in this figure.
@@
An important observation can be made by comparing Fig. 6 and Fig. 7. The gap between the SESD-based approach and the benchmarks is greater in Fig. 6 than in Fig. 7. Although an RIS with continuous phase shifts is considered in Fig. 6, the performance of the benchmark schemes that employ sub-optimal precoding vectors is very poor. When an optimal precoding design is considered in Fig. 7, the benchmarks demonstrate a better performance, although the RIS phase shifts are still sub-optimally found in these schemes. This suggests that the optimal precoder design is more important for improving performance as compared to the optimal RIS configuration design. The reason is that with the sub-optimal RIS configuration designs, only the phase shifts are affected as the reflection coefficients of the RIS are assumed to have unit modulus. On the other hand, a sub-optimal precoding design also affects the amplitude of the precoding entries, which has a large impact on the achievable sum rate.
@@ VI-C Discussion
VII Concluding Remarks and Future Outlook
@@
We studied an RIS-assisted MU-MIMO downlink communication scenario in this paper. Since the precoding is computed in the cloud in future networks, the precoding vectors must be selected from a discrete set. Furthermore, a practical RIS typically has a low phase shift resolution per element. We developed a novel framework for sum rate maximization considering precoding vectors and RIS configurations adhering to the above-mentioned limitations. We devised efficient solutions based on the SESD algorithm for both precoding and RIS configuration design and compared them against the well-known coordinate descent and nearest point designs. Numerical results demonstrated the effectiveness of the proposed designs and revealed the noticeable performance gains of the presented schemes over the commonly used benchmarks. The SESD complexity grows exponentially with respect to the problem size, making it unsuitable for practically large RIS. Hence, we developed a novel low-complexity SESD-based algorithm that divides the discrete RIS configuration optimization problem into a number of sub-problems, each solved via SESD. The effectiveness of the proposed low-complexity algorithm has also been corroborated numerically. This low-complexity approach can be utilized to solve any mixed-integer least-squares problem where the large problem size inhibits the use of conventional SD-based algorithms.
Further exploration on how to reduce the complexity of SESD-based precoding and RIS configuration optimization with minimal performance loss is important. Methods such as lattice reduction and statistical pruning as well as fixed-complexity SD algorithms are worth a closer look in the context of discrete precoding and RIS optimization. There are also many other RIS-related scenarios where the discrete phase shifts might be optimized by following the methodology outlined in this paper.